OpenStudy (anonymous):

What is the slope of a line that is perpendicular to 6x + 3y = –9?

OpenStudy (dayakar):

slope of a line ax+by+c=0 is m=-a/b in this case a=6, b=3 m1= -6/3= -2

OpenStudy (dayakar):

slope of a line which is perpendicular to above line m2 m1 * m2 =-1 (-2) * m2 =-1 m2= -1/-2 m2 = 1/2

OpenStudy (vheah):

You have to change 6x+2y=-9 to y=mx+b where m would be your slope that you can base on for your perpendicular slope. So, algebra: 1) subtract 6x from both sides 2y=-6x-9 2) divide two on both sides y=-6/2x-9/2 3) simplify y=-3x-9/2 Your slope is -3x. A perpendicular slope is the negative reciprocal of your given slope -3x which is 1/3x. 1/3x is your perpendicular slope. If you have questions on how I got to the answer, please reply! I hope I helped!